OpenStudy (anonymous):

Find the value of y when x equals 14 in this equation. -6x + 6y = -42 A. 21 B. 6 C. 7 D. -21

OpenStudy (anonymous):

plug in x=14 gives -6(14) +6y = -42 -84 +6y = -42 Add 84 to both sides gives ... 6y = 42 Divide both sides by y gives ... y = 7 .... i.e. answer C